OpenStudy (anonymous):

Write a formula for this sequence: 19, 11, 3, -5,...

OpenStudy (blues):

n - 8

OpenStudy (anonymous):

Do you mean 27-8n?

OpenStudy (blues):

If you're given that n initial = 27, yes, but just in general, n sub(+1) = n-8

OpenStudy (zarkon):

\[a_1=19\]\[a_{n+1}=a_n-8\] or \[a_n=19-8(n-1)\]